Compute the yield to maturity for a zero coupon bond with a maturity of 13 years and a face value of $1000. The bond is selling for $594.06. (Assume annual discounting.) (Round to 100th of a percent and enter as a percentage, e.g. 12.34% as 12.34)
Answer #1

Current value=Face value/(1+YTM)^time period

594.06=1000/(1+YTM)^13

(1+YTM)^13=(1000/594.06)

(1+YTM)=(1000/594.06))^(1/13)

YTM=(1000/594.06))^(1/13)-1

=4.09(Approx).